OpenStudy (anonymous):

12x + 2(5-2x)= 2x-6(7-x) I'm not sure how to solve this one at all.

OpenStudy (anonymous):

solve for x ?

OpenStudy (anonymous):

12x + 10 - 10x = 2x - 42 + 6x

OpenStudy (anonymous):

2x +10 = 8x - 42

OpenStudy (anonymous):

52 = 6x

OpenStudy (anonymous):

x = 52/6